### v2

# TRACE: Learning to Compute on Circuit Graphs

Abstract.
Learning to compute, the ability to model the functional behavior of a circuit graph, is a fundamental challenge for graph representation learning. Yet, the dominant paradigm is architecturally mismatched for this task. This flawed assumption, central to mainstream message passing neural networks (MPNNs) and their conventional Transformer-based counterparts, prevents models from capturing the position-aware, hierarchical nature of computation. To resolve this, we introduce TRACE, a new paradigm built on an architecturally sound backbone and a principled learning objective. First, TRACE employs a Hierarchical Transformer that mirrors the step-by-step flow of computation, providing a faithful architectural backbone that replaces the flawed permutation-invariant aggregation. Second, we introduce function shift learning, a novel objective that decouples the learning problem. Instead of predicting the complex global function directly, our model is trained to predict only the function shift, the discrepancy between the true global function and a simple local approximation that assumes input independence. We validate this paradigm on various circuits modalities, including Register Transfer Level graphs, And-Inverter Graphs and post-mapping netlists. Across a comprehensive suite of benchmarks, TRACE substantially outperforms all prior architectures. These results demonstrate that our architecturally-aligned backbone and decoupled learning objective form a more robust paradigm for the fundamental challenge of learning the functional behavior of a circuit graph.
conference: ; ; 00footnotetext: Equal contribution.
1. Introduction
Circuit graphs provide a fundamental abstraction for modeling computation. As directed graphs of nodes representing operations and variables, they capture the flow of computation and are crucial in various circuit modalities. Accurately modeling the computational behavior of circuit graphs is therefore a critical enabler for high-impact applications, including performance prediction (Chen et al., 2024; Xie et al., 2022; Zhang et al., 2020; Mendis et al., 2019), verification (Li et al., 2023; Selsam et al., 2018; Zhang et al., 2021), and optimization (Zuo et al., 2023; Mirhoseini et al., 2021).
This need for functional modeling has driven recent work in graph representation learning, with approaches largely divided into two families: message passing neural networks (MPNNs) and Graph Transformers. These models have been increasingly applied to capture the functionality of diverse computational graph modalities, with a significant body of work focusing on the particularly challenging domain of hardware design, including Register Transfer Level (RTL) graphs (Fang et al., 2025c, b), And-Inverter Graphs (AIGs) (Shi et al., 2023, 2024; Zheng et al., 2025; Khan et al., 2025; Wu et al., 2025; Wang et al., 2024; Liu et al., 2024), and post-mapping (PM) netlists (Shi et al., 2025b). MPNNs provide a general framework (Gilmer et al., 2017) based on a paradigm of aggregating and updating node features:
(1) $\mathbf{x}_{i}^{\prime}=\gamma\left(\mathbf{x}_{i},\,\underset{j\in\mathcal{N}(i)}{\square}\,\varphi\left(\mathbf{x}_{i},\mathbf{x}_{j},\mathbf{e}_{j,i}\right)\right),$
where $\square$ denotes a permutation-invariant aggregator (e.g., sum, mean, max) and the update function $\gamma$ and message function $\varphi$ are differentiable functions. In contrast, Transformer-based methods either flatten the graph into a sequence for global self-attention (Fang et al., 2025c, b), or incorporate graph structure via attention masks (Shi et al., 2024; Zheng et al., 2025; Fang et al., 2025a).
Figure 1. The architectural failure of MPNNs on circuit graphs. Top: The permutation-invariant aggregation in MPNNs cannot distinguish between ordered inputs (e.g., A,B vs B,A), yielding the same incorrect embedding for a position-aware operator like MUX. Bottom: Our approach processes inputs as an ordered sequence, enabling position-awareness and capturing operator-specific interactions.
Despite their success on general graph domains, we argue that these architectures are ill-suited for computation due to a fundamental mismatch between their properties and the nature of computation process in circuit graphs (as shown in Fig. 1):
  1. (1)
    MPNNs Fail to Model Input Interactions: The core architectural flaw of MPNNs is that their message functions model interactions between an operator and only a single input at a time. These independent messages are then combined by a permutation-invariant aggregator. This two-stage process makes it architecturally impossible to capture the intricate, operator-specific relationships between multiple inputs and, as a direct consequence, to model position-aware operators like the MUX (multiplexer), where input order is critical (i.e., $\texttt{MUX(S, A, B)}\neq\texttt{MUX(A, S, B)}$).
  2. (2)
    Vanilla Transformers are Hierarchy-Agnostic: Fully-connected Transformers flatten the graph into a sequence, destroying the explicit hierarchy and connectivity for modeling computation. For an expression like $y=f(g(a,b),c)$, this architecture fails to capture the intermediate dependency on $g(a,b)$ and cannot guarantee a correct computational trace.
  3. (3)
    Edge-masked Transformers Inherit MPNN Limitations: Edge-masked transformers, while respecting connectivity, often reduce to GAT-like attention mechanisms. This is functionally a weighted sum, which still inherits the fundamental limitations of the message-passing paradigm, failing to model the precise, non-linear interactions required by logical and algebraic operators.
These architectural flaws highlight a failure to model the step-by-step flow of computation. Beyond this, however, a deeper challenge lies in capturing global function that emerges from the graph’s overall topology. Even with simple components (e.g., AND and NOT gates in AIGs), a graph’s overall functionality can become highly complex due to reconvergent dependencies. For instance, consider $c=a\land b$ with $a=x\land y$ and $b=y\land z$, where $x,y,z\sim\mathcal{B}(p)$ (a Bernoulli distribution with parameter $p$). Locally, $a,b\sim\mathcal{B}(p^{2})$, and ignoring reconvergence, one would predict $c\sim\mathcal{B}(p^{4})$. However, since $y$ appears in both $a$ and $b$, $a$ and $b$ are correlated, shifting the true distribution to $c\sim\mathcal{B}(p^{3})$. This function shift, from $\mathcal{B}(p^{4})$ to $\mathcal{B}(p^{3})$, demonstrates how dependencies fundamentally alter functional behavior. Previous works that supervise directly on a node’s final global function implicitly bundle this effect into the embeddings, making it difficult for the model to distinguish true functional dependencies from spurious correlations.
Our work addresses these challenges with TRACE, a Transformer for Reasoning about Algebraic and Computational Expressions, which presents a two-fold solution. First, to resolve local architectural flaws of prior models, TRACE employs a hierarchical Transformer. Inspired by prefix notation, we represent a computation step as an ordered sequence, [operator, input_1, input_2, …], processed by a Transformer encoder with positional encoding (Figure 1). Applying this process recursively according to graph’s logical dependencies, TRACE learns a faithful, position-aware representation of each computational step. Second, to capture the global function, we introduce function shift learning, a novel objective that explicitly models the discrepancy between local and global functions, allowing the model to disentangle a node’s intrinsic behavior from the contextual effects imposed by wider graph topology.
Our experiments span a broad spectrum of circuit modalities (RTL, AIG, and PM netlists) and tasks (contrastive and predictive) across several standard benchmarks, including ITC (Corno et al., 2002), OpenCores (Albrecht, 2005), ISCAS ’89 (Brglez et al., 1989), ForgeEDA (Shi et al., 2025a), and DeepCircuitX (Li et al., 2025). The results are unequivocal: TRACE consistently and substantially outperforms prior approaches across all settings. This establishes TRACE not only as a new state of the art for circuit analysis, but as a more robust and architecturally sound paradigm for learning on circuit graphs.
2. Background
Figure 2. Illustration of a circuit graph and its computation process. This figure demonstrates the dual role of nodes within a circuit graph. Step 1 shows nodes 1 and 2 functioning as operators to compute the expressions $S\land B$ and $\lnot S$, respectively. As computation progresses to Step 2, these nodes transition to representing the intermediate variables that hold the results of these operations in step 1, which are then passed to subsequent operators for further computation.
2.1. Circuit Graphs
In our work, circuit graphs are defined as directed graphs where nodes represent either input variables or operators, and edges signify data flow. As depicted in Figure 2, this structure allows a single node to serve dual roles: it can be an operator receiving data from its source nodes and an intermediate variable whose output is consumed by other operators. Specifically, for any directed edge from a source to a target node, the target always represents an operator, while the source represents a variable—either an initial input variable or a temporary intermediate variable resulting from a previous operation. This representation effectively models the data dependencies and computational flow within the system. We focus on three types of circuit graphs from the front-end of the electronic design automation (EDA) flow: Register-Transfer Level (RTL) graphs, And-Inverter Graphs (AIGs), and Post-Mapping (PM) netlists.
2.2. Message Passing Neural Networks
Message Passing Neural Networks (MPNNs) are a dominant architectural paradigm for function learning on circuit graphs. These models can be broadly categorized into two types: synchronous and asynchronous. Synchronous MPNNs (Wu et al., 2023a; Liu et al., 2024; Wu et al., 2025; Deng et al., 2024) process all message-passing updates in parallel, a strategy designed for computational efficiency. In contrast, asynchronous MPNNs (Li et al., 2022; Shi et al., 2023, 2025b; Khan et al., 2025; Wang et al., 2024) mimic the logic simulation process by updating node representations sequentially, following a topological order, which aims to capture effective functional representations by emulating the data flow. Such asynchronous processing may introduces a potential limitation for processing circuits with significant logic depth. This limitation is often mitigated in practice, as real-world timing constraints typically restrict the maximum number of logic levels. Despite their architectural differences, both synchronous and asynchronous MPNNs rely on the conventional message passing paradigm (Gilmer et al., 2017) (Figure 1). This paradigm fundamentally struggles to capture the operator-specific and position-aware interactions among input variables.
Figure 3. Overview of our proposed framework. Left: A circuit graph, represented in both a graph view and its equivalent prefix notation, is encoded by a Hierarchical Transformer to model the computation process. Right: For predictive tasks, we introduce Function Shift Learning (FSL). Instead of directly regressing the global function, the model captures the difference between the global and local functions: $y^{FSL}=y^{global}-y^{local}$.
2.3. Graph Transformers
Despite the widespread use of MPNNs, they suffer from inherent limitations, including difficulty in capturing long-range dependencies and susceptibility to over-smoothing (Li et al., 2018) and over-squashing (Alon and Yahav, 2020). These drawbacks have motivated a shift towards Graph Transformers, which leverage global attention to address them. Existing graph transformers fall into two categories: fully-connected transformers and edge-masked transformers.
Fully-connected transformers, such as Rampášek et al. (2022); Wu et al. (2023c, b) for general graphs and Fang et al. (2025c, b) for circuit graphs, treat the graph as a flattened sequence of nodes. While this enables global self-attention, it inadvertently leads to a loss of the critical hierarchical structure and intermediate dependencies that are inherent to circuit graphs. In contrast, edge-masked transformers, including DeepGate3 (Shi et al., 2024), DeepGate4 (Zheng et al., 2025), and NetTAG (Fang et al., 2025a), integrate the graph’s topology by using the adjacency matrix to mask the attention mechanism. However, this approach often reduces the model to a Graph Attention Network (GAT)-like attention mechanism, inheriting the limitations of traditional MPNNs.
3. Method
3.1. Hierarchical Transformer
To address the limitations of MPNNs and Graph Transformers discussed in Section 1 and  2, we propose a new paradigm for encoding circuit graphs, with a position-aware Hierarchical Transformer that enables operator-specific interaction among input variables, faithfully mirroring the computational process illustrated in Figure 2.
A circuit graph $\mathcal{G}=(\mathbf{V},\mathbf{E})$, consists of primary input (PI) nodes, which have an in-degree of zero, and operator nodes, whose in-degree is determined by their type. We begin by computing the logic level of each node in topological order [^1] To handle cyclic graphs, such as sequential AIGs, we follow Khan et al. (2025) by treating all flip-flops or registers as pseudo Primary Inputs (PIs) and removing the feedback edges to compute the logical level. as follows:
(2) $level(v)=\begin{cases}0&\text{if $v$ is PI}\\ 1+\max\limits_{(u,v)\in\mathbf{E}}level(u)&\text{otherwise}\end{cases}$
The logic level defines the dataflow through the graph. Inspired by asynchronous MPNNs, we process nodes level by level. First, Primary Input (PI) nodes (level 0) are initialized with their input distribution, e.g., $\mathcal{B}(p)$, a Bernoulli distribution with parameter $p$. Then, for the nodes $v_{0}^{k},v_{1}^{k},\dots,v_{n_{k}}^{k}$ at subsequent level $k>0$, we gather their respective sets of direct predecessors, $\mathcal{N}(v_{0}^{k}),\dots,\mathcal{N}(v_{n_{k}}^{k})$, where each set is defined as $\mathcal{N}(v_{i}^{k})=\{u_{j}\in\mathbf{V}\mid(u_{j},v_{i}^{k})\in\mathbf{E}\}$.
Inspired by prefix notation, we represent each computation at level $k$ as an ordered sequence, $\mathbf{v}_{i}^{k}$, formed by placing the operator $v_{i}^{k}$ at the head, followed by its ordered input nodes from $\mathcal{N}(v_{i}^{k})$:
(3) $\mathbf{v}_{i}^{k}=[v_{i}^{k},u_{j_{1}},u_{j_{2}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}],\text{ where }u_{j_{1}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}\in\mathcal{N}(v_{i}^{k})$
To model operator-specific interactions among inputs, we apply a Transformer encoder to the sequence $\mathbf{v}_{i}^{k}$ augmented with positional encodings. The updated embedding for the operator node $v_{i}^{k}$ is taken from the Transformer’s output corresponding to the first token:
(4) $v_{i}^{k}=\text{Transformer}(\mathbf{v}_{i}^{k}+\mathbf{pos})[0],$
where $\mathbf{pos}$ represents the positional encodings. The input sequence $\mathbf{v}_{i}^{k}$ is composed of the initial embedding of the operator $v_{i}^{k}$ (e.g., a one-hot vector of its type) and the embeddings of its input nodes $\{u_{j_{m}}\}$, which have been computed in previous steps. After updating, the embedding of $v_{i}^{k}$ now represents the result of the computation at this node, i.e. an intermediate variable node.
This paradigm offers several key advantages. First, positional encodings enable position-aware aggregation for order-dependent operations, contrasting with the permutation-invariant nature of standard message passing. Second, self-attention enables rich interactions among all source nodes $u_{j_{1}},\dots,u_{j_{|\mathcal{N}(v_{i}^{k})|}}$], faithfully mirroring the computational dataflow of an operator. Finally, the sequence length depends only on in-degree: unlike general graphs with long-tailed degree distributions that incur high overhead, circuit graphs have small, tightly bounded in-degrees, preventing the memory explosion caused by quadratic complexity.
An Alternative View: Hierarchical Transformer on Prefix Notation. Our proposed method can also be interpreted as a Hierarchical Transformer operating on prefix notation, in a way that follows the logical dependencies inherent in the circuit graph. This perspective highlights how the model processes the graph by mirroring the step-by-step evaluation of an expression. A circuit graph can be converted into a prefix notation string through a pre-order traversal on its reversed edges. As illustrated in Figure 3, our proposed Hierarchical Transformer, unlike methods that simply flatten the graph into a single sequence, inherently preserves the nested, hierarchical structure of dependencies within each computation step, which allows it to align naturally with the actual flow of computation. This view also underscores the generalizability of our approach to other sequence-based problems, such as hardware model-checking, where the problem is often represented in Btor2 format (ArminBiere et al., 2018), a prefix-style representation of bit-vector formulas.
3.2. Function Shift Learning on Predictive Task
Logic-1 probability prediction is a task widely studied in prior works (Shi et al., 2023; Khan et al., 2025; Shi et al., 2024; Zheng et al., 2025; Shi et al., 2025b; Liu et al., 2024), as it serves as a key indicator of a model’s ability to capture circuit functionality. The logic-1 probability corresponds to the global function of a circuit, as defined in Definition 3.1. This function can be highly complex due to reconvergent dependencies, and computing it directly requires enumerating the joint distribution of all inputs, which incurs an exponential cost of $O(2^{k})$.
Definition 0 (Global Function).
Given an operator $\phi$, input variables $\mathbf{x}=[x_{1},x_{2},\dots,x_{k}]$ and its distribution $\mathcal{D}$, the global function is defined as $y_{\phi}^{global}=\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[\phi(x_{1},x_{2},\dots,x_{k})].$
By ignoring the dependencies among input, i.e. by assuming they are independent, we can derive the local function, which is formally stated in Definition 3.2. Although this approximation is computationally efficient with an $O(1)$ complexity, it fails to capture the true function of the circuit.
Definition 0 (Local Function).
Given an operator $\phi$, input variables $\mathbf{x}=[x_{1},x_{2},\dots,x_{k}]$ and its distribution $\mathcal{D}$, the local function is defined as $y_{\phi}^{local}=\phi\!\left(\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[x_{1}],\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[x_{2}],\dots,\mathbb{E}_{\mathbf{x}\sim\mathcal{D}}[x_{k}]\right).$
Building on these properties, we propose to learn the function shift: $y_{\phi}^{FSL}=y_{\phi}^{global}-y_{\phi}^{local},$ which measures the discrepancy between the local function (Definition 3.2) and the true global function (Definition 3.1). This formulation decouples the global function into two components: a simple local function and a function shift. Rather than predicting the complex global function directly, our model is trained to predict only the function shift. This isolates the complex contextual effects caused by reconvergence, allowing the global function to be reconstructed by simply combining the predicted shift with the local function.
Training Stage.
During the training stage, the ground-truth global and local functions for each node can be pre-computed from the training data, allowing us to determine the true function shift, $y_{i}^{FSL}$. We then train the model to regress this value, optimized with an $\mathcal{L}_{1}$ objective:
(5) $\min_{\theta}\ \mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\Big[\big|\psi(x_{i})-y_{i}^{FSL}\big|\Big],$
where $x_{i}$ is the embedding for node $i$ and $\psi(\cdot)$ is a regression head.
Algorithm 1 Inference with Function Shift
1:Circuit graph $\mathcal{G}=(V,E)$
[…]
5:for $l=1$ to $L$ do
6:  for $v\in\{u\in V:\text{level}(u)=l\}$ do
7:   $\hat{y}_{v}^{global}\leftarrow\hat{y}_{v}^{FSL}+\underbrace{\phi_{v}(\hat{y}_{u_{1}}^{global},\dots,\hat{y}_{u_{|\mathcal{N}(v)|}}^{global})}_{\text{Local Function}}$
8:Return $\hat{\mathbf{y}}^{global}$ for all nodes
Inference Stage.
At the inference stage, the true global functions are unknown. Since computing the global function of a node at any given level depends on the global functions of its predecessors from previous levels, we cannot predict them all at once. Therefore, we reconstruct the global functions iteratively, proceeding level by level through the circuit. For each node, we first compute its local function using the already-estimated global functions of its inputs. The final estimate for the node’s global function is then obtained by adding the model’s predicted function shift to this computed local function. This entire iterative process is detailed in Algorithm 1.
3.3. Contrastive Task
Contrastive learning is a standard self-supervised strategy for learning representations of circuit functionality (Wang et al., 2024; Fang et al., 2025b, c; Wu et al., 2025). The fundamental principle is to pull embeddings of functionally equivalent circuits closer together in the embedding space while pushing apart those of functionally different circuits. This process trains the model to identify and encode the discriminative features that define a circuit’s intrinsic properties, all without needing explicit labels. Following prior work, we form training instances for each circuit $\mathcal{G}$. A positive sample, $\mathcal{G}^{+}$, is created by applying a functionally equivalent transformation to $\mathcal{G}$. All other circuits within the same batch are treated as negative samples, $\mathcal{G}^{-}$. We then optimize the encoder using the InfoNCE loss (Oord et al., 2018):
(6) $\min_{\theta}\ \mathbb{E}_{\mathcal{G}\sim\mathcal{D}}\;\mathcal{L}_{\text{InfoNCE}}(\mathcal{G},\mathcal{G}^{+},\mathcal{G}^{-}).$
The objective is to structure the embedding space so that functional equivalence corresponds to proximity, mapping similar circuits to nearby points while separating from non-equivalent ones.
Table 1. Comparison of contrastive task across various modalities(%).
Model RTL AIG Netlist
Rec@1 Rec@5 Rec@10 Rec@1 Rec@5 Rec@10 Rec@1 Rec@5 Rec@10
Message Passing Neural Network
GCN 82.90 87.43 91.57 83.01 93.28 96.05 58.03 77.65 85.24
GraphSAGE 86.46 92.87 95.43 88.55 96.41 98.38 86.12 95.82 98.13
GAT 84.98 89.22 94.65 85.68 94.32 97.60 65.21 83.72 89.97
GIN 86.23 91.34 96.95 85.98 93.40 96.52 75.85 91.82 95.90
FGNN2 - - - 88.73 97.03 98.57 - - -
DeepCell - - - - - - 80.99 95.31 97.61
Graph Transformer
GraphGPS 86.94 92.13 96.37 OOM OOM OOM 45.63 61.55 69.57
SGFormer 79.45 86.57 89.50 15.43 30.88 42.19 15.83 37.06 49.73
DIFFormer 88.28 92.97 96.88 37.03 68.87 80.66 25.23 45.52 55.67
CircuitEncoder 88.27 92.97 94.52 - - - - - -
TRACE 94.45 98.74 99.89 92.68 98.65 99.51 90.81 98.48 99.44
4. Experiment
4.1. Implement Details
Dataset
In this paper, we conduct experiments on three modalities: RTL, AIG and PM netlist. For RTL, we follow prior works (Fang et al., 2025b, c) and collect data from ITC (Corno et al., 2002) and OpenCores (Albrecht, 2005). For combinational AIGs, we use ForgeEDA (Shi et al., 2025a). For Sequential AIGs, we follow DeepSeq2 (Khan et al., 2025) and extract sub-circuits from ITC (Corno et al., 2002), OpenCores (Albrecht, 2005) and ISCAS’89 (Brglez et al., 1989). For PM netlist, we follow prior work (Shi et al., 2025b) and extract sub-circuits from ForgeEDA (Shi et al., 2025a) and DeepCircuitX (Li et al., 2025). We summarize the statistics of the datasets used in both contrastive tasks (Table 2) and predictive tasks (Table 3). Since the global function in Definition 3.1 is hard to acquire, we follow previous works (Shi et al., 2023, 2025b) and approximate it using random simulations.
Table 2. Dataset for contrastive tasks with $\{min.,avg.,max.\}$.
RTL AIG PM Netlist
#Graphs 1138 67706 67728
#Nodes { 9, 102.6, 1987} {22, 162.2, 2127} {16, 88.6, 1192}
#Edges {10, 278.3, 2645 } {22, 176.0, 2275} {20, 110.5, 1361}
Depth {2, 8.9, 27} {6, 16.7, 29} {2, 5.9, 12}
Table 3. Dataset for predictive tasks with $\{min.,avg.,max.\}$.
Com. AIG Seq. AIG PM Netlist
#Graphs 9800 10007 83042
#Nodes {10, 2324.9, 45409} {23, 236.5, 1281} {12, 668.1, 4783}
#Edges {12, 3268.57, 68676} {21, 260.5, 1915} {11, 1113.9, 8914}
Depth {4, 49.1, 2657} {4, 21.0, 102} {1, 16.0, 260}
Evaluation Metrics
In this work, we evaluate our model on two types of tasks: contrastive and predictive. For the first type, the contrastive task, the goal is retrieval. Given a query circuit, the model must identify its functionally equivalent (positive) counterpart from a pool of $N$ candidate circuits. For our experiments, we set the pool size to $N=256$ for RTL and $N=1024$ for AIG and PM Netlists. We measure performance using the Recall@k (Rec@k) metric, reporting scores for $k\in\{1,5,10\}$. For the second type, the predictive tasks, we assess the model’s ability to determine node-level properties. We follow previous works (Shi et al., 2023; Khan et al., 2025; Shi et al., 2025b) and perform logic-1 probability prediction, similarity prediction and transition probability prediction. We evaluate the accuracy of these predictions using Mean Absolute Error (MAE) and the $R^{2}$ score.
4.2. Contrastive Tasks
As shown in Table 1, TRACE consistently outperforms all baselines across the RTL, AIG, and PM Netlist modalities. Traditional message-passing GNNs (GCN, GIN, etc.) and specialized encoders (CircuitEncoder, FGNN2, DeepCell) achieve moderate performance, with Rec@1 scores generally ranging from 80% to 88.7%. However, Graph Transformer variants (GraphGPS, SGFormer, DIFFormer) show highly variable results: while competitive on RTL (up to 88.28%), they suffer severe performance degradation on AIGs and PM Netlists (Rec@1 dropping as low as 15.4% or below 50%) and can face OOM errors, highlighting their difficulty in handling circuit graph hierarchy. In contrast, TRACE achieves superior and stable results, pushing Rec@1 to 94.45% (RTL), 92.68% (AIGs), and 90.81% (PM Netlists), with Rec@10 consistently to 99.9%, 99.5% and 99.4% respectively. These results, which outperforms the second-best method by a large margin on RTL, AIG, and PM Netlist modalities, demonstrates its strong generalization capacity.
Table 4. Comparison of predictive tasks on combinational and sequential AIGs.
Model Combinational AIG Sequential AIG
Logic-1 Probability Similarity Prediction Logic-1 Probability Transition Probability
R2 MAE R2 MAE R2 MAE R2 MAE
Message Passing Neural Network
GCN 0.644 0.152 0.271 0.090 0.868 0.064 0.744 0.024
GAT 0.618 0.157 0.029 0.090 0.877 0.053 0.831 0.016
GIN 0.669 0.144 0.445 0.076 0.962 0.035 0.790 0.023
GraphSAGE 0.675 0.143 0.438 0.078 0.927 0.048 0.867 0.017
DeepGate2 0.983 0.028 0.502 0.069 - - - -
PolarGate 0.493 0.192 0.021 0.113 - - - -
MGVGA 0.666 0.145 0.418 0.077 - - - -
DeepSeq2 - - - - 0.979 0.025 0.908 0.014
Graph Transformer
GraphGPS OOM OOM OOM OOM 0.971 0.026 0.901 0.012
SGFormer 0.516 0.175 -0.072 0.117 0.878 0.056 0.596 0.026
DIFFormer OOM OOM OOM OOM 0.701 0.097 0.416 0.034
DeepGate4 0.984 0.027 0.464 0.078 - - - -
TRACE 0.989 0.015 0.633 0.055 0.997 0.009 0.976 0.005
Table 5. Comparison of predictive task on PM netlists.
Metric Message Passing Neural Network Graph Transformer TRACE
GCN GraphSAGE GAT GIN DeepCell GraphGPS SGformer DIFFormer
R2 0.718 0.946 0.902 0.734 0.942 0.846 0.918 0.696 0.994
MAE 0.112 0.048 0.059 0.102 0.053 0.083 0.056 0.141 0.013
4.3. Predictive Tasks
As shown in Table 4 and Table 5, TRACE demonstrates a clear and consistent superiority across all predictive tasks by achieving minimal MAE and best $R^{2}$, validating its strong generalization. On Combinational AIGs (Table 4), where traditional message-passing baselines and Graph Transformers struggle, TRACE delivers outstanding performance. TRACE achieves an MAE of 0.015 for logic-1 probability prediction and significantly outperforms all baselines in similarity prediction with an MAE of 0.055. The performance gap is further emphasized on Sequential AIGs, whose temporal dependencies pose additional challenges. Here, TRACE achieves substantial gains, reaching a minimal MAE of 0.009 for logic-1 probability prediction and an MAE of 0.005 for transition probability prediction, demonstrating superior precision. Finally, on PM Netlists (Table 5), TRACE again surpasses all baselines, achieving a minimal MAE of 0.013. This consistent ability to minimize error across modalities, in contrast to the varied performance of other models, emphasizes the adaptability of our approach. Overall, by consistently delivering the minimal MAE scores, TRACE demonstrates its robust capacity to generalize across diverse combinational and sequential circuit graph domains.
Figure 4. Error Accumulation Analysis. We illustrate the MAE of logic-1 probability prediction across different logic levels. Levels containing fewer than 10 nodes are excluded for better visualization. Top: Results on AIGs. Bottom: Results on PM netlists.
Error Accumulation
The iterative inference procedure in Algorithm 1 raises potential concerns about error accumulation. We investigate this by analyzing performance stability across different logic levels. As shown in Figure 4, we plot the MAE for the logic-1 probability prediction task, grouping nodes by their logic level. The results demonstrate that TRACE maintains a stable and outstanding performance across all levels. This indicates the robustness of TRACE to potential error accumulation and generalizes effectively to circuit graphs of varying depths.
Table 6. Ablation study on function shift learning (FSL).
Task GraphSAGE DeepCell DeepGate4 TRACE
w/o w. w/o w. w/o w. w/o w.
AIG Prob. 0.143 0.043 - - 0.027 0.020 0.024 0.015
Sim. 0.078 0.071 - - 0.078 0.076 0.066 0.055
PM Netlist Prob. 0.048 0.046 0.053 0.047 - - 0.036 0.013
4.4. Ablation Study
We conducted an ablation study to quantify the contribution of our proposed Function Shift Learning (FSL) component. The results in Table 6, compare our model (TRACE), second-based synchronous MPNNs (GraphSAGE) and asynchronous MPNNs (DeepGate4, DeepCell) with (w.) and without (w/o) the FSL component.
The findings demonstrate that FSL is crucial for achieving high performance. Our model, TRACE with FSL, consistently outperforms its ablated counterpart across all tasks. The impact is particularly significant on the PM Netlist probability prediction, where FSL reduces the MAE from 0.036 to 0.013, a relative error reduction of 63.9%. For the AIG tasks, FSL also yields significant gains, reducing the probability prediction MAE by 37.5% (from 0.024 to 0.015) and the similarity prediction MAE by 16.7% (from 0.066 to 0.055).
While FSL provides a consistent performance boost for baseline MPNNs on AIGs, they gain only a slight boost on the PM Netlists modality. This suggests that standard MPNNs struggle to leverage FSL on circuit graphs where nodes represent more complex functions, highlighting their limitation in capturing both the global function the function shift—a gap TRACE is designed to address.
5. Conclusion
In this work, we introduced TRACE, a new paradigm for learning on circuit graphs that addresses the architectural limitations of conventional MPNNs and Transformers. By employing a novel Hierarchical Transformer and a function shift learning objective, TRACE directly models the position-aware, hierarchical nature of computation. Our extensive experiments on electronic circuits demonstrate that TRACE substantially outperforms all prior architectures, establishing a new state of the art. This work provides a proof of principle for a more architecturally sound approach to learning on circuit graphs, offering a powerful framework with potential applications across various domains.